OpenStudy (anonymous):

The product denoted by X in the followingalpha decay. 212Rn86 à AXZ + 4He2

OpenStudy (anonymous):

how would I do this?

OpenStudy (anonymous):

from the law of conservation of particles, 86=z+2 & 212=A+4 therefore,Z=84 and A=208 since the atomic number of X is 84, it is Lead

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

wow, its not so hard at all, now that I know how to solve it.